# Qualification-free convex analysis via the joint supporting subspace

### v4

@@ Abstract
In convex analysis, qualification conditions (also termed constraint qualifications) help avoid pathological behavior at domain boundaries. In this work we remove the need for such conditions by localizing to an affine subspace—the joint supporting subspace—that contains the feasible region and ensures qualification conditions hold after localization.
Our theory generalizes Borwein and Wolkowicz’s facial reduction beyond conic programs to convex programs of the form $f(x)+g(Ax)$. Intuitively, the joint supporting subspace corresponds to a bilateral facial reduction between any two convex sets. It enables simple qualification-free generalizations for a host of central results of convex analysis. These include: an exact Fenchel-Rockafellar dual; Karush-Kuhn-Tucker (KKT) optimality conditions; attained infimal convolution for convex conjugates; subdifferential sum and chain rules; and a characterization of the normal cones of the intersection of two convex sets. All generalizations reduce seamlessly to their original formulations when qualification conditions hold.
We offer a number of characterizations for the joint supporting subspace, one of which is constructive. Our proofs are self-contained, and introduce a novel theoretical framework consisting of nested normals and supporting subspaces, which simultaneously describe both the boundary of convex sets and the lattice of faces.
@@ 1 Introduction
In convex analysis, many results depend on qualification conditions, of which one prominent example found in the seminal work [rockafellarConvexAnalysis1970] by Rockafellar, is that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$,
@@
where $\operatorname{relint}$ denotes the relative interior. This condition ensures that the feasible region $\operatorname{dom}(f+g)$ is not contained in the relative boundary of either $\operatorname{dom}(f)$ or $\operatorname{dom}(g)$. When Equation 1 does not hold, the geometry of the boundary of the convex sets $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$ becomes important, and requires special treatment. Such cases are typically excluded from consideration with qualification conditions because of the added complexity, and also because such cases are often avoidable in practice.
The sum rule of subdifferentials,
$\partial(f+g)(x)=\partial f(x)+\partial g(x),$
is guaranteed by [rockafellarConvexAnalysis1970, Theorem 23.8], subject to Equation 1 (amongst other possible sufficient conditions). Other results of convex analysis with similar conditions include, e.g., the subdifferential chain rule, attainment of the infimal convolution in convex conjugates, strong duality of the Fenchel-Rockafellar dual, and Karush-Kuhn-Tucker (KKT) optimality conditions. For results concerned with optimization problems, this type of condition is usually called constraint qualifications.
@@
Qualification conditions are not merely theoretical artifacts; they are justified by counterexamples, and therefore reflect more or less accurately the true scope of applicability of the associated rules. We introduce one such counterexample for the sum rule of subdifferentials, to later show how our theory resolves this case. Take the function $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=\infty$ otherwise. Consider the indicator function $g(x)=\iota_{\mathbb{R}_{-}}(x)$, which is 0 on $\mathbb{R}_{-}$ and $\infty$ elsewhere. Naive application of the sum rule yields
$\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset.$
@@
In fact, $\partial(f+g)(0)=\partial\iota_{0}(0)=\mathbb{R}$ so the sum rule does not hold. There is no conflict with [rockafellarConvexAnalysis1970, Theorem 23.8] because qualification conditions are violated; the relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect. The standard sum rule fails in this case, though we will show that it is recovered after localization to the joint supporting subspace.
We tackle the problem of qualification conditions directly by developing geometric results describing the boundary of convex sets, allowing us to generalize a collection of results from convex analysis in one fell swoop in Section 5, with the classical formulations trivially recovered when qualifications hold (see Proposition 5.1).
The key geometric insight is as follows. For any two intersecting convex sets $C,D\subseteq\mathbb{R}^{n}$ in $\mathbb{R}^{n}$, there exists a special affine subspace $T_{a}$—the joint supporting subspace—that contains $C\cap D$ and that is uniquely determined by $C$ and $D$, i.e., $T_{a}:=T_{a}(C,D)$. We provide multiple characterizations of the joint supporting subspace in Section 4. Supporting subspaces (see Definition 6.1) are a generalization of supporting hyperplanes, first defined, to our knowledge, in [brondstedIntroductionConvexPolytopes1983, Exercise 5.4]. As intersections with supporting hyperplanes reveal exposed faces of convex sets, so do intersections with supporting subspaces reveal faces—which need not be exposed. As the name suggests, $T_{a}$ is a supporting subspace of both $C$ and $D$, which means that two faces of $C$ and $D$ are simultaneously revealed by intersection with $T_{a}$. Each of the convex sets presents a face to the other, and they meet within $T_{a}$. The utility of $T_{a}$ in convex analysis stems from two simple facts:
  1. 1.
    $C\cap D\subseteq T_{a}.$
  2. 2.
We prove these facts in Corollary 4.1 and Corollary 4.5 respectively. These two properties can be understood as follows: when reducing the ambient space to $T_{a}$, the the reduced ( localized, intersected) sets $C^{\prime}:=C\cap T_{a}$ and $D^{\prime}:=D\cap T_{a}$ preserve the original intersection ($C^{\prime}\cap D^{\prime}=C\cap D$) while satisfying the qualification condition $\operatorname{relint}(C^{\prime})\cap\operatorname{relint}(D^{\prime})\neq\emptyset$. Similarly, for convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, with $T_{a}:=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$, the localized functions $f^{\prime}:=f+\iota_{T_{a}},g^{\prime}:=g+\iota_{T_{a}}$ preserve the original sum ($f^{\prime}+g^{\prime}=f+g$) while satisfying Equation 1. Standard convex analysis results can then be applied to the localized functions $f^{\prime}$ and $g^{\prime}$, while still accurately describing the original sum $f+g$. For instance, in Corollary 5.1 we state that
$\partial(f+g)(x)=\partial f^{\prime}(x)+\partial g^{\prime}(x),$ (2)
without any condition on $f$ and $g$ other than they be proper convex functions. A host of other generalizations are obtained via minor variations on this procedure in Section 5.
This modification to the sum rule resolves the previously encountered counterexample. In that problem, where $f(x)=-\sqrt{x}$ and $g(x)=\iota_{(-\infty,0]}$, the joint supporting subspace from Definition 4.1 is $T_{a}=\{0\}$, and so
$\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x)=\partial\iota_{\{0\}}+\partial\iota_{\{0\}}=\mathbb{R},$
which indeed matches $\partial(f+g)(x)$.
1.1 Facial reduction and computational outlook
As we show in Section 5.3.1, our theory is a generalization of facial reduction, a preprocessing technique with strong theoretical guarantees in conic programming. Facial reduction reduces the problem to the smallest face of the conic constraint that contains the feasible set [massamOptimalityConditionsConeConvex1979, borweinFacialReductionConeconvex1981, borweinRegularizingAbstractConvex1981]. After reducing to this face, Slater’s condition is provably satisfied. As mentioned in [drusvyatskiyManyFacesDegeneracy2017], facial reduction leads to both a reduction in the size of the problem and improved numerical stability, and has consequently been implemented in solvers for semi-definite programs (e.g., in the MOSEK solver [permenterSolvingConicOptimization2017, fribergFacialReductionMOSEK]). The impact of facial reduction in computation inspires hope that our theory could lead to improvements in optimization algorithms in non-conic settings where constraint qualifications may fail.
Where facial reduction reduces to the minimal face of a cone, we reduce jointly to the smallest faces of two convex sets, both selected by intersection with the single joint supporting subspace. For simplicity, we frame our results as a reduction of the relevant objects (convex sets or convex functions), while keeping the ambient space $\mathbb{R}^{n}$ fixed. Our results are however naturally suited to a reduction of the ambient space to the joint supporting subspace, i.e., the joint supporting subspace would then become the new ambient space (see, e.g., Theorem 5.1). This alternative mathematical formulation would better model a reduction of the problem in a preprocessing step. Reduction of the ambient space to the joint supporting subspace in a preprocessing step may seem counterintuitive, as results in convex analysis are commonly specified only for real vector spaces, not affine spaces. This discrepancy can be mitigated by translating the affine subspace jointly with all the objects it contains until it reaches origin. But even this minor adjustment seems redundant, because, we believe, affine spaces seem to be the natural home for many results in convex analysis.
In full abstraction, an affine space is a pair $(L,U)$ consisting of a set $L$ and a real vector space $U$ (the tangent space), equipped with an addition operation $L\times U\to L$ with the property that $\forall l_{1},l_{2}\in L$ there exists a unique $u\in U$ such that $l_{1}+u=l_{2}$. For $L$ an affine subspace of $\mathbb{R}^{n}$, the tangent space is $U=L-L=\{l_{1}-l_{2}\mid l_{1},l_{2}\in L\}$. Results in convex analysis can then be generalized by distinguishing tangent vectors from points in the affine space. For example, that affine subspaces are closed under convex combinations follows from reformulating $\lambda x+(1-\lambda)y$ for $x,y\in L$ as $y+\lambda(x-y)$, by noticing that $y\in L$ and that $\lambda(x-y)$ is in the tangent space LLL-L. The tangent space appears in our definition for the separation of convex sets (Definition 4.3), and justifies our repeated use of self-differences of sets SSS-S (see, e.g., Definition 6.2, Corollary 4.1, or Definition 4.2).
1.2 Theoretical contribution
An important contribution of the present work is the theoretical framework that we build in Section 6. The central objects are nested normals (Definition 6.3) and supporting subspaces (Definition 6.2), which describe local subsets of the lattice of faces of a convex set. Their combination forms a mathematical object that is an exact dual for convex cones (see Theorem 6.2). As such, the nested normals can also be understood as a local description of the boundary of convex sets, an idea that we discuss in Section 6. With these objects, we derive results describing the intersection of convex sets, as well as the facial structure of differences of convex sets. Notable technical results include Theorem 6.3, Theorem 6.4, and Theorem 6.6. We therefore contribute a geometric framework describing the boundaries, intersections, and differences of convex sets.
@@ 1.3 Paper outline
Section 2 surveys prior works, notably facial reduction in conic programming, limiting approaches to exact subdifferential calculus, and the lexicographic characterization of the faces of convex sets. In Section 3 we introduce some notation. In Section 4 we state our results describing the joint supporting subspace, with proofs deferred to Section 7. There, we define the joint supporting subspace in Definition 4.1, characterize it with a distinct analytical formula in Theorem 4.1, and with a constructive iterative procedure in Corollary 5.7 which terminates in a number of steps no more than the ambient dimension. In Theorem 4.2 we state the main property of the joint supporting subspace: that qualification conditions hold when localizing to it. In Section 5, we state and prove generalizations of central results in convex analysis. In Section 5.3.1, we show that results of facial reduction in conic programming can be recovered from our results. In Section 6, we provide a narrative, motivated introduction to our mathematical framework, introducing key concepts and proving associated results along the way. In Section 7, we prove the results of Section 4 with the tools from Section 6.
@@ 2 Prior works
Constraint qualifications (CQs) in finite dimensions evolved from Slater’s condition [slaterLagrangeMultipliersRevisited1959], later weakened to the algebraic Mangasarian-Fromovitz CQ [mangasarianFritzJohnNecessary1967] and to the geometric Abadie CQ [abadieNonlinearProgramming1967]. The weak Slater condition, introduced by Rockafellar [rockafellarConvexAnalysis1970], uses the notion of relative interiors. Subsequent relaxed CQs include the constant rank constraint qualification condition [janinDirectionalDerivativeMarginal1984], further relaxed in [qiConstantPositiveLinear2000, andreaniRelaxedConstantPositive2012], and a CQ by [robinsonFirstOrderConditions1976], which takes the form of a stability condition now known as linear metric subregularity. We note that in his seminal work [rockafellarConvexAnalysis1970], Rockafellar used the same qualification conditions for optimization problems and subdifferential calculus, which enabled our unified treatment of optimization and subdifferential calculus.
There has been two major research directions generalizing results beyond qualification conditions. The geometric stream, to which our work belongs, focuses on the geometry of constraints and feasible sets. The sequential stream considers limits of increasingly accurate approximations, with exact results in the limit, and is related to variational analysis [mordukhovich_VariationalAnalysisApplications_2018, rockafellarVariationalAnalysis1997]. The geometric stream focuses on constraint qualifications for optimization problems, while the sequential stream emphasizes subdifferential calculus, with extensions to sequential optimality conditions. Our work extends the geometric stream to subdifferential calculus. We survey qualification-free results in finite dimensions that are comparable to our results in Section 5.
2.1 Geometric stream
In the setting of inequality-constrained convex programs, the Fritz-John condition [johnExtremumProblemsInequalities2014] is an optimality condition that holds without qualification conditions. Unlike the KKT condition, it is not a sufficient condition for optimality, and is satisfied either when KKT conditions hold, or when some degeneracy in the constraints is detected, akin to applying the KKT conditions with a suitable constraint qualification.
@@ 2.1 Geometric stream
A qualification-free necessary and sufficient condition was formulated by [ben-talCharacterizationOptimalityConvex1976], and extended to the nonsmooth case by [wolkowiczGeometryOptimalityConditions1980]. These works add a term to the KKT condition to deal with pathological cases; see [dharaOptimalityConditionsConvex2011, Chapter 6] for a treatment. Our qualification-free KKT condition in Corollary 5.6 is formulated for programs of the form $f(x)+g(Ax)$, instead of inequality-constrained convex programs, and uses faces of convex sets instead of an additional term.
@@
Facial Reduction (FR) generalized preprocessing ideas from linear programming to the setting of conic programming [massamOptimalityConditionsConeConvex1979, borweinFacialReductionConeconvex1981, borweinRegularizingAbstractConvex1981], with subsequent developments providing efficient computational methods for the reduction. An early method iteratively reduces the dimension of the problem [borweinRegularizingAbstractConvex1981, luoDualityResultsConic1997, wakiFacialReductionAlgorithms2013] until the minimal face is reached. Ideas of facial reduction can alternatively be framed with exact dual problems [borweinCharacterizationOptimalityAbstract1981], with important applications to semidefinite programming [ramanaExactDualityTheory1997, patakiStrongDualityConic2013]. For in-depth literature reviews of facial reduction and its numerous applications, we refer the reader to [tuncelStrongDualityMinimal2012, drusvyatskiyManyFacesDegeneracy2017, permenterReductionMethodsSemidefinite2017].
Many results in our work have analogues in facial reduction. For instance, the notion of a minimal face appears in different forms in Definition 4.1 and Theorem 4.1. The iterative reduction procedure of [borweinRegularizingAbstractConvex1981] finds its analogue in Theorem 7.1. We derive an exact Fenchel-Rockafellar dual in Corollary 5.5, echoing exact duals in conic programming. Finally, our qualification-free optimality condition in Corollary 5.6 generalizes similar results in the conic setting.
2.2 Sequential stream
The sequential approach to relaxing qualification conditions began with subdifferential calculus. One approach is that of $\varepsilon$-subdifferentials, which defines subgradients satisfying the subgradient inequality up to an $\varepsilon$-error, recovering the exact subdifferential in the limit of vanishing $\varepsilon$ [hiriarturruty_SubdifferentialCalculusUsing_1993, penotSubdifferentialCalculusQualification1996, mordukhovich_VariationalAnalysisApplications_2018]. A second, similar approach is that of fuzzy sum rules, which approximate the subdifferential by perturbing the point $x$ at which the subdifferential is taken, deriving exact additivity rules in the limit of vanishing perturbations [jules_FormulasSubdifferentialsSums_2002, ioffeFuzzyPrinciplesCharacterization1998].
These ideas were then applied to optimality conditions, where the condition of optimality takes the form of the existence of some sequence converging to the optimal point. One approach is to consider sequences of points converging to the point of interest [thibaultSequentialConvexSubdifferential1997], similarly to fuzzy sum rules, while another is analogous to $\varepsilon$-subdifferentials [jeyakumarInequalitySystemsGlobal1996, jeyakumarNewSequentialLagrange2003] relating them to the epigraphs of conjugate functions.
2.3 Theory of convex faces
A geometric insight that was critical to the development of our theory was the lexicographical characterization of faces [martinez-legazLexicographicalSeparationRn1987, martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_]. [martinez-legazLexicographicalSeparationRn1987] derive a separation theorem from which the characterization follows. It turned out that the lexicographical framework was not flexible enough for our needs, prompting us to develop an analogous yet independent and self-contained theory. In this paper, we recover the lexicographical characterization of convex faces in Theorem 6.4, though in different terms. Our proof is direct, and does not make use of a separation theorem, though some separation ideas are implicit in the definition of supporting subspaces (Definition 6.1).
3 Preliminaries

### v5

# Bilateral facial reduction: qualification-free subdifferential calculus and exact duality

@@ Abstract
Qualification conditions (also termed constraint qualifications) help avoid pathological behavior at domain boundaries in convex analysis. By generalizing facial reduction from conic programming to general convex programs of the form $f(x)+g(Ax)$, we provide qualification-free generalizations of several key results: an exact Fenchel-Rockafellar dual, KKT optimality conditions, an attained infimal convolution for the conjugate of a sum, subdifferential sum and chain rules, and normal cones of intersections. All our results reduce seamlessly to their original formulations when qualification conditions hold.
The core insight is that for a sum of two convex functions, there is an affine subspace—the joint supporting subspace— that contains the feasible region, and such that qualification conditions hold when restricting the effective domain of each function to it. We offer a number of characterizations for the joint supporting subspace, including one that obtains the affine subspace via iterative, bilateral reduction between the two domains. In our proofs, which are self-contained, we develop a structured induction on faces where inductive steps are associated with normal vectors nested in supporting subspaces (a generalization of supporting hyperplanes). With this tool, we characterize the facial structure of the difference of two convex sets from the facial structures of the individual convex sets.
@@ 1 Introduction
In convex analysis, many results require so-called qualification conditions (sometimes called constraint qualifications), such as
@@
This condition can be understood as “avoiding the boundary”; it ensures that the feasible region $\operatorname{dom}(f+g)$ is not contained in the relative boundary of either $\operatorname{dom}(f)$ or $\operatorname{dom}(g)$. When Equation 1 fails to hold, the geometry of the boundary of the convex sets $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$ starts to matter, and such cases may require special treatment.
@@
Qualification conditions are not merely theoretical artifacts; they are justified by counterexamples, and therefore reflect more or less accurately the true scope of applicability of the associated formulas. We introduce one such counterexample for the sum rule of subdifferentials. (We will later show how our theory resolves this case.) Take the function $f(x):=-\sqrt{x}$ when $x\geq 0$ and $f(x):=\infty$ otherwise. Consider the indicator function $g(x)=\iota_{(-\infty,0]}(x)$, which is 0 on $(-\infty,0]$ and $\infty$ elsewhere. Naive application of the sum rule yields
$\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset,$
@@
when in fact, $\partial(f+g)(0)=\partial\iota_{\{0\}}(0)=\mathbb{R}$, so the sum rule does not hold in this case. There is no contradiction in this failure because the qualification condition in Equation 1 is not satisfied in this case (otherwise, the sum rule would be guaranteed by [23, Theorem 23.8]). Indeed, the relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect.
Our approach generalizes facial reduction from the conic setting to general convex programs of the form f(x)+g(Ax)f(x)+g(Ax)—let us here consider the special case $f(x)+g(x)$. We identify a single affine subspace $T_{a}\subseteq\mathbb{R}^{n}$ that simultaneously regularizes both functions and contains the feasible region. The joint supporting subspace $T_{a}$ is defined for any two convex sets (see Definition 4.1), which are usually the domains of two convex functions. We can then reduce the effective domains of each function to $T_{a}$: $f^{\prime}:=f+\iota_{T_{a}},g^{\prime}:=g+\iota_{T_{a}}$. For these reduced functions, qualification conditions are then guaranteed to hold, while preserving the original problem in the sense that $f+g\equiv f^{\prime}+g^{\prime}$. With these reduced functions we modify the sum rule to resolve our previously encountered counterexample, by using the reduced functions on the r.h.s: $\partial(f+g)(x)=\partial f^{\prime}(x)+\partial g^{\prime}(x)$. In our counter-example, the joint supporting subspace is $T_{a}=\{0\}$, and the modified sum rule yields
$\partial f^{\prime}(0)+\partial g^{\prime}(0)=\partial\iota_{\{0\}}(0)+\partial\iota_{\{0\}}(0)=\mathbb{R},$
which indeed matches $\partial(f+g)(0)$. We note that the reduced functions $f^{\prime}$, $g^{\prime}$ are trivial in this example only because the feasible region is itself trivial. In general, our reduction will preserve the functions’ behaviour inside—and to some extent outside—the feasible region.
In this manner, in Section 5 we generalize a collection of results from convex analysis, with the classical formulations trivially recovered when qualifications hold (see Proposition 5.1). Specifically, we find a modified Fenchel-Rockafellar dual for which strong duality always holds, problem-independent Karush-Kuhn-Tucker (KKT) optimality conditions, always-attained infimal convolution for the convex conjugates of sums of pairs of functions, qualification-free subdifferential sum and chain rules, and a characterization of the normal cones of the intersection of two convex sets.
We provide a number of characterizations for $T_{a}$ in Section 4. We specify the joint supporting subspace $T_{a}$ in terms of the difference of domains $\operatorname{dom}(f)-\operatorname{dom}(g)$, where the set difference is $C-D:=\{c-d\mid c\in C,d\in D\}$. Other characterizations relate $T_{a}$ to minimal faces of the domains containing the feasible region (Theorem 4.1, Corollary 4.2, and Corollary 4.3). One of our characterizations (Theorem 7.1) consists of an iterative reduction procedure which recovers $T_{a}$ and terminates in a number of steps no more than the ambient dimension. This reduction can be performed from any point in the feasible region, using so-called nested normal cones (see Definition 6.3). We show that the reduction at a point does not in fact depend on the choice of the point, which means that the reduction need only be performed once for a given optimization problem. It is known that facial reduction in the conic setting leads to a reduction of the size of the problem and improved numerical stability [7], which gives reason to believe that our generalization may have similar benefits.
A contribution of the present work is in our proof technique, found in Section 6. The central objects are nested normals (Definition 6.3) and supporting subspaces (Definition 6.2), which together describe the local lattice of faces of a convex set. We establish a structured induction via nested normals, which traverses all supporting subspaces, with induction steps associated with nested normals (see Corollary 6.1). With nested normals and our structural induction, we characterize the minimal supporting subspace containing a set (see Theorem 6.3). We also characterize the facial structure of a difference of convex sets CDC-D at 0 exclusively from the local facial structure of the individual convex sets $C$ and $D$ at $C\cap D$ (see Corollary 6.2). These results enable us to prove our characterizations of the joint supporting subspace that are presented in Section 4.
@@ 1.1 Paper outline
Section 2 surveys prior works, notably facial reduction in conic programming, limiting approaches to exact subdifferential calculus, and the lexicographic characterization of the faces of convex sets. In Section 3 we introduce some notation. In Section 4 we state analytic and constructive characterizations of the joint supporting subspace, with proofs deferred to Section 7. In Theorem 4.2 we state the main property of the joint supporting subspace: that qualification conditions hold when localizing to it. In Section 5, we state and prove generalizations of central results in convex analysis. In Section 6, we build the core theoretical framework. In Section 7, we prove the results of Section 4 with the tools of Section 6.
@@ 2 Prior works
Constraint qualifications (CQs) in optimization started with Slater’s condition [24] and much work has since been aimed at finding relaxations of these conditions [15, 1, 22]. Notably, a qualification condition using relative interiors (the weak Slater) was introduced by [23], who used the same qualification conditions for optimization problems and subdifferential calculus, which enabled our unified treatment.
There are a number of qualification-free alternatives to classical results, which come at the cost of some added complexity. One approach is to consider approximations of the core objects, with qualification-free exact results obtained in the limit of increasingly accurate approximation. This type of result includes $\varepsilon$-subdifferentials [8], and fuzzy sum rules [9, 11]. Similarly, sequential optimality conditions have been formulated [25, 10]. Sequential approaches are related to the broader theory of variational analysis.
Another kind of qualification-free result is obtained by considering the geometry of domain boundaries, with the notable example of facial reduction in conic optimization. Our work is the first, to our knowledge, to extend this geometric methodology to qualification-free subdifferential calculus.
@@
In this vein, a qualification-free necessary and sufficient condition was formulated by [2], and extended to the nonsmooth case by [27], which achieve qualification-free characterization of minimizers by adding a term to the KKT conditions. Our qualification-free KKT condition in Corollary 5.6 is formulated for programs of the form $f(x)+g(Ax)$, instead of inequality-constrained convex programs, and uses a reduction instead of an additional term.
@@
Our theory extends Facial Reduction (FR) from conic programs to programs of the form $f(x)+g(Ax)$. The field of facial reduction started by generalizing preprocessing ideas from linear programming to the setting of conic programming [18, 3, 5], with subsequent developments providing efficient computational methods. One computational method consists in iteratively reducing the dimension of the problem [5, 14, 26] until the minimal face is reached, an approach that we generalize in Theorem 7.1. Ideas of facial reduction can alternatively be framed as exact dual problems [4], with important applications to semidefinite programming [21, 20].
From convex geometry, a line of work that enabled the development of our theory was the lexicographical characterization of faces [17, 16]. It turned out that the lexicographical framework was not flexible enough for our needs, prompting us to develop an analogous self-contained theory, recovering the lexicographical characterization of convex faces in Theorem 6.2 in different terms. Specifically, our framework based on nested normals allows us to specify a special kind of structural induction on supporting subspaces (and thereby faces) in Corollary 6.1. This induction is powerful because each inductive step is associated with a single nested normal. Additionally, the use of nested normals allows us to specify a reduction procedure that uses nested normal cones at each step, instead of individual nested normals (see Corollary 4.4 and Theorem 7.1). This formulation provides structural insight and allows, in most cases, termination in fewer steps by dropping many dimensions at once in some steps.
Shortly after the first version of our work was made public, an independent paper by [12] was made available which specifies a reduction for convex programs of the form $f(x)+g(x)$. Their work has an algorithmic focus: they show feasibility detection and fast reductions with extended duals, by defining a notion of niceness for the domains of the convex functions. In contrast, our work is more geometric, providing analytical characterizations of the joint supporting subspace in Section 4 and qualification-free results stated with this object in Section 5, including a reduction for programs in the slightly more general form $f(x)+g(Ax)$, and subdifferential calculus sum and chain rules. Additionally, the proof techniques of the two papers seem remarkably distinct: they lift to cones, whereas we introduce nested normals in the original space.
3 Notation and Preliminaries